Solve the system by substitution. Show all work!
2x – 2y = 2
y = 3x – 17

Answer by SHUgrad05(58)

2x – 2y = 2
y = 3x – 17
2x-2(3x-17)=2
2x-6x+34=2
-4x+34=2
-4x=-32
x=8
2(8)-2y=2
16-2y=2
-2y=-14
y=7
{8,7}